# Enhancing Target-unspecific Tasks through a Features Matrix

### v2

@@ 4.1 Benchmark Setting
Table 3: Domain generalization.
@@
Compared Methods. In three generalization experiments, we compare our method with CLIP (ICML2021) (Radford et al., 2021), CoOp (IJCV2022) (Zhou et al., 2022b), CoCoOp (CVPR2022) (Zhou et al., 2022a), MaPLe (M) (CVPR2023) (Khattak et al., 2023a), PromptSRC (S) (ICCV2023) (Khattak et al., 2023b), FAP (NIPS2024) (Zhou et al., 2024) and existing DePT (CVPR2024) (Zhang et al., 2024). The DePT is an easy-to-use method. We process frozen CLIP using a linear probe method. Comparison methods are based on ViT-B/16 architecture for fair comparison.
@@
Implementation Details. We employ a CLIP model based on the ViT-B/16 architecture. For the PromptSRC-based and MaPLe-based frameworks, we set the visual and textual embedding length to 4. We set the easy-to-use module $\gamma$ to $0.1$ and matching scores (top and low) $\beta$ to $5$. Training for 30 epochs for a base-to-novel setting in the first 9 transformer layers. Training for 20 epochs for domain generalization setting and cross-dataset evaluation setting in the first 3 transformer layers. We use an SGD optimizer with a learning rate of 0.0025 on a single GPU. Following the hand-crafted CLIP, we use the hand-crafted template sets from (Radford et al., 2021).
4.2 Datasets
In Table 5, the datasets cover multiple recognition tasks. These datasets are all open-source datasets. We conducted base-to-novel generalization experiments, few-shot learning experiments, and cross-dataset generalization experiments on 11 datasets. We conduct domain generalization experiments on four variant datasets.
@@ 4.2 Datasets
Table 5: Training and testing datasets. dataset consists of 11 image classification datasets and four variant datasets of ImageNet.
4.3 Base-to-Novel Generalization Task
In the base-to-novel generalization task, the datasets are divided into base and novel classes. The model is trained on the base classes in a 16-shot setting, and tested on both the base and novel classes across 11 different datasets. The number of classes for base and novel is the same, which means that all classes in a dataset are evenly divided into two groups of classes. The process of dividing all classes in the dataset is randomly selected. HM refers to harmonic mean. This experiment evaluates the generalization ability of methods. In Table 2, our method demonstrates significant improvements on all $11$ datasets on HM. Our method significantly enhances the performance of base classes while also improving the accuracy of zero-shot CLIP on novel classes. In terms of average performance, our method based on PromptSRC (s) attains $85.70\%$ accuracy on the base classes, and $77.35\%$ accuracy on the novel classes. Our method, as an easy-to-use module, achieves performance improvements for various baseline frameworks, such as CoOp, CoCoOp, MaPLe (M), and PromptSRC (S).
4.4 Cross-Dataset Generalization Task
We train our model with 16 shots on ImageNet and test on 10 other unseen datasets. This experiment aims to validate the potential of our method in a wide range of cross-dataset transfers. Our method for training ImageNet is lower than that of DePT. As shown in Table 4, our method based on PromptSRC shows competitive performance in $/10$ 10 over the generic and flexible method DePT. These findings suggest that our method excels in achieving better generalization across a diverse range of unseen datasets.
4.5 Domain Generalization Task
We train our model using the ImageNet in 16 shots and test its performance on 4 different variants of the ImageNet. In Table 3, our method based on PromptSRC demonstrates the highest average improvement of $1.27\%$ over the PromptSRC method. Additionally, when compared to DePT, our method consistently outperforms it across all ImageNet variant datasets. This indicates that our method is designed to enhance domain generalization.
5 Further Studies
5.1 Learning Depth and Embeddings Length
In Table 6, we note that increasing the learning depth generally increases the performance based on PromptSRC. As the number of layers increases to $11$, the HM value decreases. It indicates excessive fine-tuning of the model, causing it to lose CLIP generalization. In Table 7, our findings indicate that the performance based on PromptSRC reaches its peak when the length of embeddings is set to $4$ on HM for an average of 11 datasets. Our ablation studies are based on keeping all other settings unchanged.
Table 6: Analysis of learning depth based on PromptSRC for an average of 11 datasets.
Learning Depth 1 3 5 7 9 11
HM 77.11 78.03 79.12 80.01 81.32 80.21
Table 7: Analysis of embeddings length based on PromptSRC for an average of 11 datasets.
Prompt Length 1 2 4 6 8 10
HM 77.01 80.05 81.32 80.30 79.11 78.54
5.2 Analysis of $\gamma$ for $\mathcal{L}_{\text{CL}}$
In Table 8, it is observed that as the number of $\gamma$ to 0.1, there is a peak in HM for an average of 11 datasets. We conduct analysis based on the PromptSRC model.
5.3 Analysis of Top-$\beta$ and Low-$\beta$
In Table 9, it is observed that as the number of Top-$\beta$ and Low-$\beta$ to 5, there is a peak in HM ($81.32\%$). We conduct analysis based on the PromptSRC model for an average of 11 datasets. We keep other hyperparameters unchanged.
Average (11 datasets)
0.05 0.1 0.5 0.9
HM ($\gamma$) 80.13 81.32 81.00 79.88
Average (11 datasets)
3 4 5 6
HM ($\beta$) 78.21 79.81 81.32 80.51
5.4 Applying to Other ViT Instances
Table 10: Applying to other ViT instances.
Average (11 datasets)
ViT-B/32 ViT-B/16 ViT-L/14
HM (ViT) 79.70 81.32 84.20
5.5 Inference Stage Computational Cost
In Table 11, the compute cost analysis is performed using the SUN397 dataset over 10 epochs on a single GPU. While our method (row 3) may have a slower inference speed as a result of multiple cosine similarity calculations. The number of parameters we introduced has remained relatively consistent and stable in training. This implies that usually, our parameter quantity is lower than that of DePT.
@@ 6.1 Vision Language Models
Recently, researchers have demonstrated the strong generalization capability of Visual Language Models (VLMs) (Alayrac et al., 2022) which involve training on large-scale datasets of image-text pairs. Such as CLIP (Radford et al., 2021), which is a prominent and straightforward framework among existing VLMs. The strong generalization capability of CLIP has made it a foundation for many methods in adapting pre-trained VLMs for downstream tasks (Sanghi et al., 2022; Maaz et al., 2022; Bangalath et al., 2022; Zhang et al., 2021). To enhance the generalization ability of VLMs, researchers have explored various approaches. Some approaches involve enhancing the text encoder or the visual encoder (Vaswani et al., 2017). By improving the capabilities of the text encoder, the model can better capture the semantics and contextual information in the textual input. Similarly, enhancing the visual encoder allows the model to extract more informative and discriminative features from the visual input. Another approaches deeply fuse the visual and text knowledge within the model. Using a large number of images for training is indeed one of the most effective methods.
6.2 Prompt Tuning in Vision Language Models
@@
Prompt tuning (Lester et al., 2021) is a commonly employed technique in the field of Natural Language Processing (NLP) for training on downstream tasks. Leveraging text prompts, which are instructions given to the language model component of VLMs, is a prevalent practice to improve task comprehension. Full fine-tuning and linear probes are two commonly employed approaches for adapting VLMs to downstream tasks. However, the constraints of both methods have prompted research into innovative techniques influenced by prompt tuning within the realm of Natural Language Processing. CoOp (Zhou et al., 2022b) and CoCoOp (Zhou et al., 2022a) fine-tuning the CLIP model for few-shot image recognition by optimizing a continuous set of embeddings within the textual branch. The image-conditional prompt utilized in CoCoOp significantly contributes to improving generalization to unseen classes. Moreover, some approaches (Yao et al., 2023b, a) constrain the proposed learnable prompts contain the essential general knowledge and prior distribution learning. By conditioning prompts on visual features, CoCoOp (Zhou et al., 2022a) ensures that the language model focuses on pertinent visual information when making predictions. In addition to prompt tuning, approach (Zhang et al., 2024) introduces a generic and flexible approach to align its vision and language representations.
7 Limitations and Conclusion
@@
Due to our method incorporating multiple contrastive learning processes, the training speed is slower. Moreover, we plan to investigate the potential of our method in other downstream tasks (Sanghi et al., 2022).
@@
In this paper, we propose a Features Matrix (FM) for vision-language models. Our method incorporates multiple hand-crafted prompts to extract general information as a pre-trained features matrix from CLIP to enhance generalization. The matrix of pre-trained features can delve into the semantics of different hand-crafted prompts at a more profound level, which reduces the risk of forgetting the essential general knowledge of pre-trained CLIP. Importantly, our method is compatible with representative prompt learning frameworks for textual or multi-modal prompts as a generic and flexible module.
@@
This paper presents work whose goal is to advance the field of Machine Learning. None of these points we feel must be specifically highlighted here.

### v4

@@ 4 Main Generalization Tasks
We evaluate our method on generalization tasks. As evidenced by systematic benchmarking in Figure 1, our framework, when synergistically integrated with MaPLe (Khattak et al., 2023a) and PromptSRC (Khattak et al., 2023b), achieves state-of-the-art performance across three critical generalization dimensions (base-to-novel generalization, domain generalization, and cross-dataset generalization) spanning 11 heterogeneous benchmarks. Notably, our method surpasses the existing easy-to-use DePT (Zhang et al., 2024) by a significant margin in generalization scenarios through low-resource training.
@@ 4.1 Benchmark Setting
Table 3: Domain generalization. We train our model using the ImageNet in 16 shots and test its performance on 4 different variants of the ImageNet. When compared to DePT, our method consistently outperforms it across all ImageNet variant datasets.
@@
Compared Methods. In three generalization experiments, we compare our method with CoOp (IJCV2022) (Zhou et al., 2022b), CoCoOp (CVPR2022) (Zhou et al., 2022a), MaPLe (CVPR2023) (Khattak et al., 2023a), PromptSRC (ICCV2023) (Khattak et al., 2023b), FAP (NIPS2024) (Zhou et al., 2024), and DePT (CVPR2024) (Zhang et al., 2024). The DePT is an easy-to-use method. We process frozen CLIP using a linear probe method. Comparison methods are based on the ViT-B/16 architecture for fair comparison.
@@
Implementation Details. We employ CLIP (Radford et al., 2021) model based on the ViT-B/16 architecture. For the PromptSRC-based and MaPLe-based frameworks, we set the visual and textual embedding length to 4. We set the easy-to-use module $\gamma$ to $0.1$ and matching scores (top and low) $\beta$ to $5$. Training for 30 epochs for a base-to-novel setting in the first 9 transformer layers. Training for 20 epochs for the domain generalization setting and the cross-dataset evaluation setting in the first 3 transformer layers. We use a SGD optimizer with a learning rate of 0.0025 on a single GPU. Following the hand-crafted CLIP, we use the hand-crafted template sets from (Radford et al., 2021).
Datasets. In Table 10, the datasets cover multiple recognition tasks including ImageNet (Deng et al., 2009) and Caltech101 (Fei-Fei et al., 2004) which consists of generic objects, OxfordPets (Parkhi et al., 2012), StanfordCars (Krause et al., 2013), Flowers102 (Nilsback & Zisserman, 2008), Food101 (Bossard et al., 2014), and FGVCAircraft  (Maji et al., 2013) for fine-grained classification, SUN397 (Xiao et al., 2010) for scene recognition, UCF101 (Soomro et al., 2012) for action recognition, DTD (Cimpoi et al., 2014) for texture classification, and EuroSAT  (Helber et al., 2019) which consists of satellite images. We leverage datasets such as ImageNetA (Hendrycks et al., 2021b), ImageNet-R (Hendrycks et al., 2021a), ImageNet-Sketch (Wang et al., 2019), and ImageNetV2 (Recht et al., 2019) to assess the model’s performance across different domain distributions.
4.2 Base-to-Novel Generalization Task
In the base-to-novel generalization task, the datasets are divided into base and novel classes. The model is trained on the base classes in a 16-shot setting, and tested on both the base and novel classes across 11 different datasets. The number of classes for base and novel is the same, which means that all classes in a dataset are evenly divided into two groups of classes. The process of dividing all classes in the dataset is randomly selected. HM refers to harmonic mean. The HM evaluates the generalizable and non-generalizable ability of our methods.
Discussion. In Table 2, our method demonstrates significant improvements on all $11$ datasets on HM. We found that our performance is higher than DePT (Zhang et al., 2024) (CVPR2024) when integrated with different methods. Compared to FAP (Zhou et al., 2024) (NIPS2024), our method significantly enhances the performance of base classes while also improving the accuracy on novel classes. In terms of average performance, our method based on PromptSRC (S) attains $85.70\%$ accuracy on the base classes, and $77.35\%$ accuracy on the novel classes. Our method, as an easy-to-use module, achieves performance improvements for various baseline frameworks, such as CoOp, CoCoOp, MaPLe (M), and PromptSRC (S).
4.3 Domain Generalization Task
We train our model using the ImageNet in 16 shots and test its performance on 4 different variants of the ImageNet. The training ImageNet is a non-generalizable setting in Table 3 (column-1). This experiment aims to validate the potential of our method in domain shifts (Fang et al., 2023).
Discussion. In Table 3, our method based on PromptSRC demonstrates the highest average improvement of $1.27\%$ over the PromptSRC method. Additionally, when compared to DePT, our method consistently outperforms it across all ImageNet variant datasets.
4.4 Cross-Dataset Generalization Task
We train our model with 16 shots on the ImageNet dataset and test the model on 10 other unseen datasets. The training ImageNet is a non-generalizable setting in Table 4 (column-1). This experiment aims to validate the potential of our method in a wide range of cross-dataset transfers.
Discussion. As shown in Table 4, our method based on PromptSRC shows competitive performance in $/10$ 10 over the generic and flexible method DePT. These findings suggest that our method excels in achieving better generalization across a diverse range of unseen datasets. However, our method for training ImageNet is lower than that of DePT.
5 Further Studies
In this section, we discuss further studies, focusing on the impact of embedding length and depth, computational cost, applying to other ViT instances, analysis of $\gamma$ for $\mathcal{L}_{\text{CL}}$, and analysis of Top-$\beta$ and Low-$\beta$ for pre-trained features matrix.
5.1 Learning Depth
In Table 5, we note that increasing the learning depth generally increases the performance based on PromptSRC. As the number of layers increases to $11$, the HM value decreases.
5.2 Embeddings Length
In Table 7, our findings indicate that the performance based on PromptSRC reaches its peak when the length of embeddings is set to $4$ on HM for an average of 11 datasets. Our ablation studies are based on keeping all other settings unchanged. It indicates excessive fine-tuning of the model, causing it to lose CLIP generalization.
5.3 Analysis of $\gamma$ for $\mathcal{L}_{\text{CL}}$
In Table 9, it is observed that as the number of $\gamma$ increases to 0.1, there is a peak in HM for an average of 11 datasets. We conduct analysis based on the PromptSRC model.
5.4 Analysis of Top-$\beta$ and Low-$\beta$
In Table 8, it is observed that as the number of Top-$\beta$ and Low-$\beta$ increases to 5, there is a peak in HM. We conduct analysis based on the PromptSRC model for an average of 11 datasets. We keep other hyper-parameters unchanged.
Table 5: Analysis of learning depth based on PromptSRC for an average of 11 datasets. HM refers to harmonic mean.
Learning Depth 1 3 5 7 9 11
HM 77.11 78.03 79.12 80.01 81.32 80.21
Table 6: Applying to other ViT instances.
Avg. (11 datasets)
B/32 B/16 L/14
HM (ViT) 79.70 81.32 84.20
Table 7: Analysis of embeddings length based on PromptSRC for an average of 11 datasets. HM refers to harmonic mean.
Prompt Length 1 2 4 6 8 10
HM 77.01 80.05 81.32 80.30 79.11 78.54
Avg. (11 datasets)
3 4 5 6
HM ($\beta$) 78.21 79.81 81.32 80.51
5.5 Applying to Other ViT Instances
Avg. (11 datasets)
0.05 0.1 0.5 0.9
HM ($\gamma$) 80.13 81.32 81.00 79.88
5.6 Inference Stage Computational Cost
In Table 11, the compute cost analysis is performed using the SUN397 dataset over 10 epochs on a single GPU. Our method (row 3) may have a slower inference speed as a result of multiple cosine similarity calculations. The number of parameters we introduced has remained relatively consistent and stable in training. This implies that usually, our parameter quantity is lower than that of DePT.
@@ 5.6 Inference Stage Computational Cost
Table 10: Training and testing datasets. The dataset consists of 11 image classification datasets and four variant datasets of ImageNet.
@@ 6 Related Work
Vision-Language Models (VLMs). Recently, researchers have demonstrated strong generalization capability of Vision-Language Models (VLMs) (Alayrac et al., 2022; Shao et al., 2025), which involve training on large-scale datasets of image-text pairs. Using a large number of samples for training is indeed one of the most effective methods (Yang et al., 2021; Wang et al., 2023). Such as CLIP (Radford et al., 2021), which is a prominent and straightforward framework among existing VLMs. The strong generalization capability of CLIP has made it a foundation for many methods in adapting pre-trained VLMs for downstream tasks (Sanghi et al., 2022; Maaz et al., 2022; Bangalath et al., 2022; Zhang et al., 2021; Wang et al., 2022; Xu et al., 2025). To enhance the generalization ability of VLMs, researchers have explored various approaches (Zhou et al., 2025b, a). Some approaches involve enhancing the text encoder or the visual encoder (Vaswani et al., 2017). By improving the capabilities of the text encoder, the model can better capture the semantics and contextual information in the textual input. Similarly, enhancing the visual encoder allows the model to extract more informative and discriminative features from the visual input.
@@
Prompt Tuning in VLMs. Prompt tuning (Zhang & Tian, 2025; Pan et al., 2024; Cao et al., 2025) is a commonly employed technique in the field of Natural Language Processing (NLP) for training on downstream tasks (Yu et al., 2025, 2024b, 2024a). Leveraging text prompts, which are instructions given to the language model component of VLMs, is a prevalent practice to improve task comprehension (Cui et al., 2025b). Full fine-tuning and linear probes are two commonly employed approaches for adapting VLMs to downstream tasks. The constraints of both methods have prompted research into innovative techniques influenced by prompt tuning within the realm of VLMs. CoOp (Zhou et al., 2022b) and CoCoOp (Zhou et al., 2022a) fine-tune the CLIP model for few-shot image recognition by optimizing a continuous set of embeddings within the textual branch. The image-conditional prompt utilized in CoCoOp significantly contributes to improving generalization to unseen classes. By conditioning prompts on visual features, CoCoOp (Zhou et al., 2022a) ensures that the language model focuses on pertinent visual information when making predictions. Moreover, some approaches (Yao et al., 2023b, a; Cui et al., 2024, 2025a) constrain the learnable prompts to contain the essential general knowledge. In addition, the approach (Zhang et al., 2024) introduces a generic and flexible approach to align its vision and language representations.
7 Future Works and Limitations
@@
In the future, we plan to investigate the potential of our method in other tasks (Tian et al., 2024; Yang et al., 2023; Li et al., 2019; Yu et al., 2023; Li et al., 2024c; Zhang et al., 2022; Wang et al., 2021; Li et al., 2025a, c; Jia et al., 2025; Liu et al., 2023b, a; Li et al., 2024b) and downstream scenarios (Pan et al., 2025; Jin et al., 2025; Li et al., 2025b; Liu et al., 2024; Wu et al., 2024). Due to multiple processes of our method, the training speed is slower.
8 Conclusion
@@
Prompt learning is a promising method for adapting pre-trained Visual-Language Models (VLMs) for target-specific classification tasks. However, these optimization frameworks exhibit limited efficacy when applied to target-unspecific or generalizable scenarios. This performance degradation may stem from overfitting-induced catastrophic forgetting, where the model loses generalizable knowledge critical for adapting to unseen tasks during the fine-tuning process. In this paper, we propose a Features Matrix (FM) for vision-language models. Our method incorporates multiple hand-crafted prompts to extract general information as a pre-trained features matrix from CLIP to enhance generalization. The matrix of pre-trained features can delve into the semantics of different hand-crafted prompts at a more profound level, which reduces the risk of forgetting the essential general knowledge of pre-trained CLIP. Our method focuses on the challenge of various target-unspecific tasks and scenarios across a wide range of real datasets. Importantly, our method is compatible with representative prompt learning frameworks for textual or multi-modal prompts as a generic and flexible module.
Acknowledgement
XM T was supported in part by the NSFC No. 62222117. J Y was supported in part by the NSFC No. 62125201 and U24B20174.
@@ Acknowledgement
The societal implications of our work involve democratizing the availability of potent AI resources, as our plug-and-play method can achieve impressive performance even in the absence of extensive labeled data.